The floor of the recreation room in a house is a retangle. The sum of its length, width, and height is 42 ft. The sum of its length and height is double its width. The area of its floor is 280 square feet. What are the dimensions of the recreation room?
PLEASE HELP!

Answer by checkley71(8403)

L*W=280
L+W+H=42
L+H=2W
THUS
L=280/W
280/W+H=2W
280W-2W=-H
278W=-H
H=-278W
THEN
280W+W-278W=42
281W-278W=42
3W=42
W=14
L=280/14
L=20
20+14+H=42
H=42-20-14
H=42-34
H=8
PROOF
20*14=280
280=280
L=20,W=14,H=8